The chickens at Colonel Thompson's Ranch have a mean weight of 1550 g, with a standard deviation of 200 g. The weights of the chickens are closely approximated by a normal curve. Find the percent of all chickens having weights more than 1370 g. Use the standard normal table.

It is expected that ____ percent of all chickens have weights of more than 1370 g. (Type an integer or decimal rounded to two decimal places as needed.)
Answer #1

Solution :

Given that ,

mean = = 1550

standard deviation = = 200

P(x > 1370) = 1 - P(x < 1370)

= 1 - P[(x - ) / < (1370 - 1550) / 200)

= 1 - P(z < -0.9)

= 1 - 0.1841

= 0.8159

81.59% of all chickens have weights of more than 1370